From nch@roe.ac.uk Fri Aug 25 12:03:56 2006 Date: Fri, 25 Aug 2006 11:15:06 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Johann Bryant , Mike Read , Nigel Hambly , Nicholas Cross , Bob Mann , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , John Taylor , Jim Emerson , Malcolm Stewart , Mike Irwin , Mark Holliman , Peredur Williams , Stephen Warren Subject: WFAU VDFS Science Archive meeting minutes, 25 Aug 2006 Minutes of WFAU VDFS Science Archive meeting: 25th August 2006 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, PMW, RSC, MAR, NJC, JB, MSH Apologies: JPE, AL, RGM, ETWS, JMS DONM: 10am, Friday 1st September 2006 in the Vista Hut meeting room Actions discharged this week: ----------------------------- ACTION: JB to review Hardware/OS/DBMS design doc to note areas that need updating and itemise new sections required Discharged ACTION: NCH to remove udsSource/udsDetection and associated tables from the EDR, pointing users to DR1 or Nottingham EDR on those web pages Discharged ACTION: NCH to communicate with JPE over possible VDFS overview docs based on existing docs VDF-WFA-WSA-001 and 005. Discharged; JPE agreed that this was a good idea and has the old docs to hack a new overview. ACTION: JB to buzz the NeSC UKLighters to see what's going on Discharged; no reply as yet to phone or email. ACTION: NJC to import the UKIRT Faint Standards in the archive. Discharged; scripts have been written. The team has decided to keep these data within the confines of the main DB rather then import as an external database. NCH will review the scripts and do the importation. ACTION: JB to run CU19 (release DB preparation) for the Transit survey. Discharged; DB has been prepared on the load server but will be copied over to the public server for testing through the interface. Actions partly discharged but continuing: ----------------------------------------- ACTION: ETWS to finalise the browser pages for the two survey-ish non-surveys. Continues but nearly done. The following from last week partly done but continue: p) Release calibration and transit survey data as non-survey databases - ACTION: MAR, NCH & JB l) Review global archive schema to remove extraneous objects - ACTION: ALL, as we go along. NCH: VDF-WFA-VSA-002 Science Requirements Analysis PMW: " " " 003 Management and Planning NCH: " " " 004 Interface Control JB: " " " 006 Hardware/OS/DBMS Design NJC: " " " 007 Database Design MAR: " " " 008 User Interface RSC: " " " 009 Software Architecture Design, incl. local data flow RGM: " " " 010 VO Enablement ACTION: ALL to draft their respective documents (see above) by the 1st September. Continue; Drafts of 002 and 004 are on the modified development web page; others will appear over the next week. Actions carried forward from 18/08/06 meeting: ---------------------------------------------- The following from last week continue: b) Merge in the development CVS branch to the main - on hold until RSC fully tested development software - ACTION: RSC d) Debug CU7 for deprecated frame sets and deletions on rerun - ACTION: NCH i) Enhance seaming to use quality bit information - ACTION: RSC (on hold until after h) k) Ensure ingest code can cope with 06A missing data quirks and any associated new attributes - ACTION: ETWS m) Insert Millenium Galaxy Catalogue as external survey; rearrange external catalogues onto batch catalogue server - ACTION: NJC & JB o) Add in a default row for every detector appearing in every detection table (for schema consistency when querying merged sources and individual detections) - ACTION: RSC & NCH - all except b) are on hold until after the documentation for the review has been prepared. RSC anticipates doing b) within the next week or so, and notes that existing funcionality in the Main branch should (!) be unaffected. Specific points and new actions: -------------------------------- Project management: NCH noted that the next VDMT will occur within the next week or so, date TBD (1st or 5th September). NCH noted that of the 9 Vista PSPs currently under consideration and review, so far 2 have replied to the request for user requirements input (VVV and VMC). These responses (and any others that arrive in time) will be folded in to the requirements analysis document for the review. WFCAM update: Nothing new to report this week Comments and issues arising from CASU fortnightly minutes: No new minutes as of 25 Aug AM. Networking: NCH noted that 06A data is now appearing as OK_TO_COPY on /data/apm19_c/. JB noted that we are still awaiting a script to update the 3 buggy reprocessed nights in 05A; NCH emphasised the need to try to get transfer and ingest tasks running in the background, despite the current major push on review documentation. WSA Operations: All quiet on the operations front. Hardware: No news this week. Software: NCH noted that he has reorganised the archive development homepage (http://www.roe.ac.uk/~nch/wfcam) in advance of the review so that document drafts can be scrutinised by any interested parties from there. No other news this week owing to work on documentation. Survey Data Release: Nothing to report this week. Non-survey Data Release: Nothing new this week. Astrogrid deployment: MSH updated the team on the state of play as regards DSA authentication and security. Still some issues to be sorted out by the AG development team. Miscellaneous: MAR reported that he has written a short article on the WSA for the UKIRT Newsletter. Finally, NCH noted that AL and JPE have both been enquiring about the IPAC Montage package (http://montage.ipac.caltech.edu/), wondering if this may save us some work on some of the larger scale VO imaging applications. MAR volunteered to look into this in a spare moment.